# Learning to Act without Actions

### v1

@@ 6.3 Inspecting the Latent Action Space
To better understand what LAPO learns, we generate a UMAP projection (McInnes & Healy, 2018) of the latent action space for each of the 16 Procgen environments, shown in Figure 5. For most games, the structure of the latent action space is highly interpretable, with distinct clusters of latent actions closely corresponding to the true discrete actions. We also observe that the structure of the latent space varies greatly across environments. In BigFish, Chaser, Dodgeball, Maze, Miner, Plunder, and StarPilot, latent actions aggregate together in well-defined clusters aligned with the true actions. In other environments, including Climber, CoinRun, FruitBot, Jumper and Ninja, the latent action space exhibits more fragmented structure. This split roughly corresponds to environments with higher or lower degrees of partial observability. In the latter group of environments, the pixel observation shows only part of the level, cropped locally to the player character. Thus, when the player is moving, most observations will show parts of the level that were hidden in previous observations. In order for the FDM to accurately predict the next observation, the IDM needs to encode this off-screen information into the latent action, which we hypothesize leads to more complex structures in the latent space. However, as demonstrated in Section 6.1, the latent-space policy still performs well downstream even on environments with greater degrees of partial observability. Moreso, we find that vector quantization has a large impact in simplifying the structure of the latent action space. We generate the same UMAP projection for an ablation of our method without vector quantization in Appendix A.2. We note that for several environments, including those with and without partial observability, latent action clusters corresponding to distinct true actions are at times duplicated. For example, in BigFish, each true action sees four latent action clusters—likely reflecting how taking the same true action in different parts of the state-space can have differing effects. These results strongly support our hypothesis that vector quantization leads to simpler latent representations with less fragmentation within actions.
@@ 7 Conclusion
Our results thus suggest that LAPO can serve as a useful approach for obtaining rapidly-adaptable pretrained policies from web-scale action-free demonstration data, e.g. in the form of videos. We believe LAPO could serve as an important stepping stone to unlocking the web-scale unsupervised pretraining paradigm that has proven effective in other domains like language and vision (Brown et al., 2020; Radford et al., 2021; Ramesh et al., 2021a). Toward this goal, we are excited about future work aimed at scaling up the dynamics models and decoder to more powerful architectures that can consider multiple timesteps and richer observations. Such models may enable efficient learning of generalist latent action policies in complex multi-task settings, all within a single LAPO instance, while enabling more effective generalization to tasks unseen during training.
@@ A.1 Offline decoding results
Figure 6: Test performance of the latent policy combined with a latent action decoder that is trained on an action-labelled offline dataset of a certain size, consisting of $(z_{t},a_{t})$ tuples. An effective decoder can be trained with only a few hundred samples although performance generally plateaus before reaching the performance of our proposed online PPO fine-tuning approach.
@@ A.3 Continuous control experiments
Figure 8: Mean test returns over the course of training for PPO fine-tuning of a pretrained LAPO policy compared to PPO from scratch (averaged across 2 seeds) on continuous control tasks. These experiments used the same hyperparameters for training the latent IDM and for behavior cloning as in Procgen experiments.
A.4 ILPO reproduction
Figure 9: Reproduction of Figure 5.a from Edwards et al. (2019) for a single level from coinrun. When applied to data from the full distribution of levels, rather than just a single level, the ILPO FDM consistently collapsed in terms of which discrete latent achieves the minimum of $\mathcal{L}_{\text{min}}$, leading to collapse of the policy too. Results here are not exactly comparable to results from Edwards et al. (2019) since it is unknown which specific coinrun level was used by the authors.
A.5 Hyperparameters

### v2

@@ 6.3 Inspecting the Latent Action Space
To better understand what LAPO learns, we generate a UMAP projection (McInnes et al., 2018) of the latent action space for each of the 16 Procgen environments, shown in Figure 5. For most games, the structure of the latent action space is highly interpretable, with distinct clusters of latent actions closely corresponding to the true discrete actions. We also observe that the structure of the latent space varies greatly across environments. In BigFish, Chaser, Dodgeball, Maze, Miner, Plunder, and StarPilot, latent actions aggregate together in well-defined clusters aligned with the true actions. In other environments, including Climber, CoinRun, FruitBot, Jumper and Ninja, the latent action space exhibits more fragmented structure. This split roughly corresponds to environments with higher or lower degrees of partial observability. In the latter group of environments, the pixel observation shows only part of the level, cropped locally to the player character. Thus, when the player is moving, most observations will show parts of the level that were hidden in previous observations. In order for the FDM to accurately predict the next observation, the IDM needs to encode this off-screen information into the latent action, which we hypothesize leads to more complex structures in the latent space. However, as demonstrated in Section 6.1, the latent-space policy still performs well downstream even on environments with greater degrees of partial observability. Moreso, we find that vector quantization has a large impact in simplifying the structure of the latent action space. We generate the same UMAP projection for an ablation of our method without vector quantization in Appendix A.2. We note that for several environments, including those with and without partial observability, latent action clusters corresponding to distinct true actions are at times duplicated. For example, in BigFish, each true action sees four latent action clusters—likely reflecting how taking the same true action in different parts of the state-space can have differing effects. These results strongly support our hypothesis that vector quantization leads to simpler latent representations with less fragmentation within actions.
@@ 7 Conclusion
Our results thus suggest that LAPO can serve as a useful approach for obtaining rapidly-adaptable pretrained policies from web-scale action-free demonstration data, e.g. in the form of videos. We believe LAPO could serve as an important step in unlocking the web-scale unsupervised pretraining paradigm that has proven effective in other domains like language and vision (Brown et al., 2020; Radford et al., 2021; Ramesh et al., 2021a). Toward this goal, we are excited about future work aimed at scaling up the world model and policy to more powerful architectures that can consider multiple timesteps and richer observations. Such models may enable efficient learning of generalist latent action policies in complex, multi-task settings, all within a single LAPO instance, while enabling more effective generalization to tasks unseen during training.
@@ A.1 Offline decoding results
Figure 6: Test performance of the latent policy combined with a latent action decoder that is trained on an action-labelled offline dataset of a certain size, consisting of $(z_{t},a_{t})$ tuples. An effective decoder can be trained with only a few hundred samples although performance generally plateaus before reaching the performance of our proposed online (RL) decoding approach.
@@ A.3 Continuous control experiments
Figure 8: Mean test returns over the course of training for online decoding of a LAPO latent policy compared to PPO from scratch (averaged across 2 seeds) on continuous control tasks. These experiments used the same hyperparameters for training the latent IDM and for behavior cloning as in Procgen experiments.
A.4 Hyperparameters
@@
A.5 ILPO reproduction
Figure 9: Reproduction of Figure 5.a from Edwards et al. (2019) for a single level from coinrun. When applied to data from the full distribution of levels, rather than just a single level, the ILPO FDM consistently collapsed in terms of which discrete latent achieves the minimum of $\mathcal{L}_{\text{min}}$, leading to collapse of the policy too. Results here are not exactly comparable to results from Edwards et al. (2019) since it is unknown to us which specific coinrun level was used by the authors.